OpenStudy (anonymous):

What is the solution to the equation? a + (–0.4) = –3 a =

OpenStudy (lolacole12):

Let's solve your equation step-by-step. a−0.4=−3a Step 1: Add 3a to both sides. a−0.4+3a=−3a+3a 4a−0.4=0 Step 2: Add 0.4 to both sides. 4a−0.4+0.4=0+0.4 4a=0.4 Step 3: Divide both sides by 4. 4a 4 = 0.4 4 a=0.1 Answer: a=0.1

OpenStudy (mathmale):

Too bad you've chosen that user name. Next time, try one that's more positive (e. g., Avidlearner). You need to find a such that the equation is true. a + (–0.4) = –3 a =

OpenStudy (anonymous):

@Juliusmattos12 can you help

OpenStudy (mathmale):

First, rewrite (+)(-) as ONE sign. Ditch the parentheses.

OpenStudy (lolacole12):

A=0.1

OpenStudy (anonymous):

oh ya like mathmale :(

OpenStudy (lolacole12):

Ok bye

OpenStudy (mathmale):

Ya, lilke mathmale :) Do the necessary work to change the (+)(-) to just one sign

OpenStudy (mathmale):

a + (–0.4) = –3 a =

OpenStudy (mathmale):

Hint: +(-0.4) = ??

OpenStudy (skullpatrol):

Add 0.4 to each side of the equation.

OpenStudy (anonymous):

not what i mean that is a bad name

OpenStudy (mathmale):

a + (–0.4) = –3 become a - 0.4 = -3. Please follow skullpatrol's suggestion now.

OpenStudy (skullpatrol):

$$\huge a + (-0.4) + 0.4 = -3 + 0.4$$